What is the density of the water when:

a. The weight of an empty graduated cylinder was 50 g.

b. 20 mL of water was placed in the graduated cylinder.

c. The weight of the graduated cylinder and water was 70 g.


1
Expert's answer
2021-11-22T03:00:02-0500

V= 20 ml

Mass of water m = 70 g - 50 g = 20 g

Density "= \frac{m}{V} = \frac{20}{20} = 1 \; g/ml"